# TimeMachine: A Time Series is Worth 4 Mambas for Long-term Forecasting

### v1

@@ 4.2 Experimental Environment
Baseline Models: We compared our model, TimeMachine, with 11 SOTA models, including iTransformer [21], PatchTST [23], DLinear [33], RLinear [18], Autoformer [29], Crossformer [35], TiDE [6], Scinet [19], TimesNet [30], FEDformer [37], and Stationary [20]. Although another variant of SSMs, namely S4 [12], exists, we do not include it in our comparison because TiDE [6] has already demonstrated superior performance over S4.
@@ 4.3 Quantitative Results
We demonstrate TimeMachine’s performance in supervised long-term forecasting tasks in Table 2. Following the protocol used in iTransformer [21], we set all baselines fixed with $L=96$ and $T=\{96,192,336,720\}$, including our method. For all results achieved by our model, we utilized the training-related values mentioned in Section 4. In addition to the training hyperparameters, we set default values for all Mambas: Dimension factor $D=256$, local convolutional width $=2$, and state expand factor $N=1$. We provide an experimental justification for these parameters in Section 5. Table 2 clearly shows that our method demonstrates superior performance compared to all the strong baselines in almost all datasets. Moreover, iTransformer [21] has significantly better performance than other baselines on the Traffic and Electricity datasets, which contain a large number of channels. Our method also demonstrates comparable or superior performance on these two datasets, outperforming the existing strong baselines by a large margin. This demonstrates the effectiveness of our method in handling LTSF tasks with varying number of channels and datasets.
In addition to Table 2, we conducted experiments with TimeMachine using different look-back windows $L=\{192,336,720\}$. Table 3 and Supplementary Table 2 demonstrate TimeMachine’s performance under these settings. An examination of these tables reveals that the implementation of extended look-back windows markedly enhances the performance of our method across the majority of the datasets examined. This also demonstrates TimeMachine’s capability for handling longer look-back windows while maintaining consistent performance.
@@
Figure 2: Average performance (in MSE) comparison among TimeMachine and the latest SOTA baselines with input sequence length $L=96$. The center of the circle represents the maximum possible error. The closer to the boundary, the better the performance.
@@
Figure 3: Qualitative comparison between TimeMachine and second-best-performing methods from Table 2. Observations are demonstrated from the test set for the case of $L=96$ and $T=720$ with a randomly selected channel and window frame of $100$ time points.
Following iTransformer [21], Figure 2 demonstrates the normalized percentage gain of TimeMachine with respect to three other SOTA methods, indicating a clear improvement over the strong baselines. In addition to general performance comparison using MSE and MAE metrics, we also compare the memory footprints and scalability of our method against other baselines in Figure 4. We measured the GPU memory utilization of our method and compared it against other baselines, with results for other baseline models taken from the iTransformer [21] paper. To ensure a fair comparison, we set the experimental settings for our method similar to those of iTransformer.
@@ 5.3 Ablation of Residual Connections
Studies have shown the effectiveness of residual connection, including models using SSMs [1] and CNNs [14]. In this section, we justify the two residual connections in our architecture: one from $E_{2}$ to the output of the two inner Mambas, and the other from $E_{1}$ to the output of $P_{1}$. These residual connections help stabilize training and reduce overfitting, especially for the smaller datasets with channel independence. Supplementary Table 1 provides experimental justification, where the Res. column indicates the presence (✓) or absence (✗) of residual connections. We observe clear improvement on both datasets when residual connections are used. This motivated us to include residual connections in our architecture, and all results presented in Tables 2 and 3 incorporate these connections.
@@ 5.4 Effects of Mambas’ Local Convolutional Width
Table 4: Ablation experiment on the local convolution width with input sequence length $L=96$.
@@ 5.6 Ablation on Mamba Dimension Expansion Factor
We also experimented with the dimension expansion factor ($E$) of the Mambas, demonstrated in Supplementary Figure 3. Increasing the block expansion factor does not lead to consistent improvements in performance. Instead, higher expansion factors come with a heavy cost in memory and training time. Therefore, we set this value to 1 by default in all Mambas and report the results in Tables 2 and  3.
@@ 6 Conclusion
This research is supported in part by the NSF under Grant IIS 2327113 and the NIH under Grants R21AG070909, P30AG072946, and R01HD101508-01.
We would thank the University of Kentucky Center for Computational Sciences and Information Technology Services Research Computing for their support and use of the Lipscomb Compute Cluster and associated research computing resources.
@@
Figure 1: Performance (MSE) comparison concerning a diverse range of dropouts with input sequence length, $L=96$.
@@
Figure 2: Qualitative comparison between TimeMachine and second best-performing methods from Table  2. Observations are demonstrated from the test set for the case of $L=96$ and $T=720$ with a randomly selected channel and window frame of $100$ time points.

### v2

@@ 4.2 Experimental Environment
Baseline Models: We compared our model, TimeMachine, with 11 SOTA models, including iTransformer [21], PatchTST [23], DLinear [34], RLinear [18], Autoformer [29], Crossformer [36], TiDE [6], Scinet [19], TimesNet [30], FEDformer [38], and Stationary [20]. Although another variant of SSMs, namely S4 [12], exists, we did not include it in our comparison because TiDE [6] has already demonstrated superior performance over S4. Similarly, as Flowformer [31] is not as competitive as iTransformer and other Transformer-based models, following [21], we did not include it.
@@ 4.3 Quantitative Results
We demonstrate TimeMachine’s performance in supervised long-term forecasting tasks in Table 2. Following the protocol used in iTransformer [21], we set all baselines fixed with $L=96$ and $T=\{96,192,336,720\}$, including our method. For all results achieved by our model, we utilized the training-related values mentioned in Subsection 4.2. In addition to the training hyperparameters, we set default values for all Mambas: Dimension factor $D=256$, local convolutional width $=2$, and state expand factor $N=1$. We provide an experimental justification for these parameters in Section 5. Table 2 clearly shows that our method demonstrates superior performance compared to all the strong baselines in almost all datasets. Moreover, iTransformer [21] has significantly better performance than other baselines on the Traffic and Electricity datasets, which contain a large number of channels. Our method also demonstrates comparable or superior performance on these two datasets, outperforming the existing strong baselines by a large margin. This demonstrates the effectiveness of our method in handling LTSF tasks with varying number of channels and datasets.
In addition to Table 2, we conducted experiments with TimeMachine using different look-back windows $L=\{192,336,720\}$. Table 3 and Supplementary Table 2, demonstrate TimeMachine’s performance under these settings. An examination of these tables reveals that the implementation of extended look-back windows markedly enhances the performance of our method across the majority of the datasets examined. This also demonstrates TimeMachine’s capability for handling longer look-back windows while maintaining consistent performance.
@@
Figure 2: Average MSE comparison of TimeMachine and SOTA baselines with $L=96$. The circle center represents the maximum possible error. Closer to the boundary indicates better performance.
@@
Figure 3: Qualitative comparison between TimeMachine and the second-best method (Table 2) on a test set example with L=96, T=720, and a randomly selected channel (best viewed zoomed-in).
Following iTransformer [21], Figure 2 demonstrates the normalized percentage gain of TimeMachine with respect to three other SOTA methods, indicating a clear improvement over the strong baselines. In addition to the general performance comparison using MSE and MAE metrics, we also compare the memory footprints and scalability of our method against other baselines in Figure 4. We measured the GPU memory utilization of our method and compared it against other baselines, with their results taken from the iTransformer [21] paper. To ensure a fair comparison, we also included Flowformer [31] and vanilla Transformer [28], and set the experimental settings for our method similar to those of iTransformer.
@@ 5.3 Ablation of Residual Connections
Studies have shown the effectiveness of residual connection, including models using SSMs [1] and CNNs [14]. In this section, we justify the two residual connections in our architecture: one from $E_{2}$ to the output of the two inner Mambas, and the other from $E_{1}$ to the output of $P_{1}$. Both of them use element-wise additions and help stabilize training and reduce overfitting, especially for the smaller datasets with channel independence. Supplementary Table 1 provides experimental justification, where the Res. column indicates the presence (✓) or absence (✗) of residual connections. We observe clear improvement on both datasets when residual connections are used. This motivated us to include residual connections in our architecture, and all results presented in Tables 2 and 3 incorporate these connections.
@@ 5.4 Effects of Mambas’ Local Convolutional Width
Table 4: Ablation results on the local convolution width with $L=96$.
@@ 5.6 Ablation on Mamba Dimension Expansion Factor
We also experimented with the dimension expansion factor ($E$) of the Mambas, which is used to expand the input dimension, with results shown in Supplementary Figure 3. Increasing the block expansion factor does not lead to consistent improvements in performance. Instead, higher expansion factors come with a heavy cost in memory and training time. Therefore, we set this value to 1 by default in all Mambas and report the results in Tables 2 and  3.
In addition to these sensitivity analyses, we also demonstrated performance comparison between 1 and 2 levels in Supplementary Table 3. Considering a balance between performance and memory footprint, we used two levels.
6 Strengths and Limitations
TimeMachine outperforms numerous baselines, including transformer-based methods, across benchmark datasets and additionally demonstrates memory efficiency and stable performance across varying look-back and prediction lengths. Unlike transformer-based methods that have quadratic complexity, our method has linear complexity. While TimeMachine achieves top-ranked performance in most cases, it ranks second on the Weather dataset with small $T$, highlighting an area for future improvement. Moreover, as shown in Figure 3, there is potential for enhancing alignment with ground truth.
@@ 7 Conclusion
This research is supported in part by the NSF under Grants 2327113 and 2433190 and the NIH under Grants R21AG070909, P30AG072946, and R01HD101508-01. We would like to thank the University of Kentucky Center for Computational Sciences and Information Technology Services Research Computing for their support and use of the Lipscomb Compute Cluster and associated research computing resources.
@@
Figure 1: Performance (MSE) comparison concerning a diverse range of dropouts with input sequence length $L=96$.
@@
Figure 2: Qualitative comparison between TimeMachine and second-best-performing methods from Table  2. Visualization is provided for the test set with $L=96$ and $T=720$ with a randomly selected channel and a window frame of $100$ time points.
@@
Table 3: Ablation experiment on the level of TimeMachine with input sequence length $L=96$ and $T=\{96,192,336,720\}$.
Prediction ($T$$\rightarrow$ 96 192 336 720
$\mathcal{D}$ Level MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 1 0.367 0.393 0.420 0.418 0.437 0.424 0.460 0.455
2 0.364 0.387 0.415 0.416 0.429 0.421 0.458 0.453
Electricity 1 0.143 0.237 0.164 0.256 0.175 0.271 0.212 0.303
2 0.142 0.236 0.158 0.250 0.172 0.268 0.207 0.298